# Bias Evaluation and Mitigation in Retrieval-Augmented Medical Question-Answering Systems

### v2

@@ Introduction
Medical question-answering (QA) systems powered by large language models (LLMs) have shown remarkable progress in knowledge-intensive tasks, promising valuable clinical decision support 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15. Despite their advancements, LLMs often suffer from issues like factual inaccuracies and hallucinations, particularly critical in high-stakes domains like healthcare. Retrieval-Augmented Generation (RAG) addresses these limitations by integrating external knowledge bases to enhance factual accuracy and minimize hallucinations 16, 17, 18, 19. Specifically, RAG-based approaches retrieve relevant evidence, thus significantly improving response reliability.
@@
However, introducing external retrieval also brings new challenges. Recent studies indicate that biases—often associated with sensitive attributes such as race, gender, or socioeconomic status—can persist or even be exacerbated within both retrieval and generation stages of RAG systems, potentially compromising fairness and reliability 20, 21.
@@
Although prior research has investigated biases within end-to-end generative models, less attention has been directed toward disparities arising specifically during the RAG pipeline 22. In this context, disparities can emerge when retrieval mechanisms systematically neglect or inadequately represent particular groups characterized by sensitive attributes such as race and gender. Moreover, generation models conditioned on prompts containing sensitive identity cues (e.g., “This African American patient”) risk reinforcing stereotypes or neglecting clinically relevant nuances. Given the growing emphasis on fairness and equity in healthcare AI, understanding how sensitive identity attributes influence retrieval and generation outcomes becomes crucial.
@@
To address this critical gap, we propose a unified framework to systematically evaluate and mitigate biases in medical RAG pipelines. Specifically, we first generate query variants that explicitly incorporate sensitive attributes (e.g., race: Caucasian, African American, Asian, Hispanic; gender: male, female, non-binary) to uncover disparities in retrieval outcomes and QA accuracy. Second, we design and compare several bias mitigation strategies, including Chain-of-Thought (COT) filtering 23 that encourages step-by-step reasoning, Counterfactual filtering that verifies consistency across varied identity contexts, and Adversarial prompts designed to prevent over-reliance on identity terms. Additionally, we introduce a Majority Vote mechanism 24, aggregating model outputs across multiple attribute-specific variants, further improving robustness and fairness.
@@
We validate our approach across multiple diverse benchmarks—MedQA25, MedMCQA26, MMLU27, and the open-ended EquityMedQA22. By leveraging fairness metrics such as Demographic Parity (DP)28 and Equal Opportunity (EO)29, our experimental results quantitatively demonstrate significant biases within existing retrieval and generation processes and highlight how our proposed mitigation methods substantially reduce these disparities.
@@
Recent advancements in large language models (LLMs) and Retrieval-Augmented Generation (RAG) have shown significant promise in medical question-answering systems. However, ensuring the fairness and trustworthiness of these systems, especially in high-stakes medical contexts, remains a critical challenge. Ni et al. 21 provide a comprehensive survey outlining key aspects of trustworthy RAG systems, highlighting reliability, privacy, safety, fairness, explainability, and accountability as critical dimensions of trust. They identify that fairness in RAG systems requires special attention in both the retrieval and generation stages, as biases introduced during retrieval can propagate to generation, potentially exacerbating disparities.
Several studies further explore bias in retrieval-augmented systems, specifically within healthcare contexts. Levra et al.20 emphasized the risks of demographic biases that can be inadvertently amplified through retrieval processes in medical QA systems. Similarly, Pfohl et al. 22 introduced EquityMedQA, an open-ended dataset explicitly designed to test demographic biases in medical QA models, demonstrating disparities arising due to sensitive attributes such as race and gender. Recent methods have proposed sophisticated bias mitigation strategies to address these fairness issues. Chain-of-Thought (COT) prompting 23, 30 encourages explicit reasoning steps, potentially reducing reliance on demographic stereotypes. Counterfactual filtering 31 alters sensitive attributes to check model consistency and minimize discriminatory outcomes. Additionally, adversarial filtering and Majority Vote aggregation 32 have effectively reduced biases and promoted more equitable model outcomes by aggregating diverse demographic perspectives. Despite these advancements, challenges persist. Retrieval methods still sometimes select outdated or irrelevant information, as evidenced by error cases involving intersectional demographics (e.g., Asian non-binary individuals), identifying a gap in robustness and generalization capabilities. Ongoing work thus emphasizes the importance of developing trustworthy RAG systems, with comprehensive frameworks proposed to address reliability, privacy, and fairness comprehensively 21.
@@ Methods
We select four representative general-purpose large language models (LLMs) for integration into our Retrieval-Augmented Generation (RAG) framework, including DeepSeek-R1-8B , DeepSeek-R1-70B33, Meta-Llama-3-8B34, and PMC-LLaMA-13B 35. These models were chosen based on their proven capabilities in general reasoning, instruction-following tasks, and suitability for adaptation into domain-specific contexts, such as medical question-answering.
We apply four strategies to mitigate demographic bias in Retrieval-Augmented Generation (RAG). First, a Plain (baseline) condition involves no explicit intervention. Second, Chain-of-Thought (COT) Filtering 23, 30 guides models to produce step-by-step reasoning that isolates medical facts from demographic descriptions, aiming to lessen undesired influences of attributes like race or gender. Third, Counterfactual Filtering 31 prompts the model with methodically altered demographic labels and checks for inconsistencies or discriminatory behaviors in the resulting answers. Lastly, Adversarial Prompt Filtering reformulates queries to minimize reliance on socially sensitive markers, thus preventing the model from overfitting to or amplifying potential biases. We compare each filtering method’s output regarding accuracy, retrieval patterns, and fairness measures during inference.
@@
To further combat biases that may persist for a single demographic instance, we incorporate a majority voting approach 32 across multiple variants of the same question. Concretely, for each question in MedMCQA, MedQA, MMLU, and EquityMedQA, we generate distinct demographic variants by substituting race (Caucasian, African American, Asian, Hispanic) and gender (male, female, non-binary) into the query stem. For example, an original question “Which medication is recommended for a patient with chest pain?” can yield up to 12 variants if we combine four race attributes and three gender attributes. In practice, if both race and gender are not always simultaneously varied, we produce 4 to 6 variants (e.g., only race or only gender) depending on the experiment.
@@
(2) ROUGE-L. For open-ended QA such as EquityMedQA 22, which lacks fixed answer options, we measure correctness by comparing the generated text to a reference solution using ROUGE-L. This metric quantifies the overlap of the longest common subsequence (LCS) between the model’s output and the reference. A higher ROUGE-L indicates greater alignment with the reference’s content, helping detect factual completeness in a free-text generation.
@@
We adopt four corpora in our retrieval pipeline, each chunked into short snippets: (1) PubMed 36 for biomedical abstracts,(2) Medical Textbooks 25 for domain-specific knowledge, and (3) Wikipedia 37 for more general context. In addition, we combine these sources into a larger MedCorp if cross-domain retrieval is desired. Each snippet is indexed and retrieved via different retriever types, including a lexical approach (BM25) 38 and semantic encoders (Contriever) 39, 40. By default, we retrieve $k=15$ snippets per query; if multiple retrievers are used, we employ Reciprocal Rank Fusion (RRF) 41 to merge results.

### v3

@@ Introduction
Medical question-answering (QA) systems powered by large language models (LLMs) have shown remarkable progress in knowledge-intensive tasks, promising valuable clinical decision support 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16. Despite their advancements, LLMs often suffer from issues like factual inaccuracies and hallucinations, particularly critical in high-stakes domains like healthcare. Retrieval-Augmented Generation (RAG) addresses these limitations by integrating external knowledge bases to enhance factual accuracy and minimize hallucinations 17, 18, 19, 20. Specifically, RAG-based approaches retrieve relevant evidence, thus significantly improving response reliability.
@@
However, introducing external retrieval also brings new challenges. Recent studies indicate that biases—often associated with sensitive attributes such as race, gender, or socioeconomic status—can persist or even be exacerbated within both retrieval and generation stages of RAG systems, potentially compromising fairness and reliability 21, 22.
@@
Although prior research has investigated biases within end-to-end generative models, less attention has been directed toward disparities arising specifically during the RAG pipeline 23. In this context, disparities can emerge when retrieval mechanisms systematically neglect or inadequately represent particular groups characterized by sensitive attributes such as race and gender. Moreover, generation models conditioned on prompts containing sensitive identity cues (e.g., “This African American patient”) risk reinforcing stereotypes or neglecting clinically relevant nuances. Given the growing emphasis on fairness and equity in healthcare AI, understanding how sensitive identity attributes influence retrieval and generation outcomes becomes crucial.
@@
To address this critical gap, we propose a unified framework to systematically evaluate and mitigate biases in medical RAG pipelines. Specifically, we first generate query variants that explicitly incorporate sensitive attributes (e.g., race: Caucasian, African American, Asian, Hispanic; gender: male, female, non-binary) to uncover disparities in retrieval outcomes and QA accuracy. Second, we design and compare several bias mitigation strategies, including Chain-of-Thought (COT) filtering 24 that encourages step-by-step reasoning, Counterfactual filtering that verifies consistency across varied identity contexts, and Adversarial prompts designed to prevent over-reliance on identity terms. Additionally, we introduce a Majority Vote mechanism 25, aggregating model outputs across multiple attribute-specific variants, further improving robustness and fairness.
@@
We validate our approach across multiple diverse benchmarks—MedQA26, MedMCQA27, MMLU28, and the open-ended EquityMedQA23. By leveraging fairness metrics such as Demographic Parity (DP)29 and Equal Opportunity (EO)30, our experimental results quantitatively demonstrate significant biases within existing retrieval and generation processes and highlight how our proposed mitigation methods substantially reduce these disparities.
@@
Recent advancements in large language models (LLMs) and Retrieval-Augmented Generation (RAG) have shown significant promise in medical question-answering systems. However, ensuring the fairness and trustworthiness of these systems, especially in high-stakes medical contexts, remains a critical challenge. Ni et al. 22 provide a comprehensive survey outlining key aspects of trustworthy RAG systems, highlighting reliability, privacy, safety, fairness, explainability, and accountability as critical dimensions of trust. They identify that fairness in RAG systems requires special attention in both the retrieval and generation stages, as biases introduced during retrieval can propagate to generation, potentially exacerbating disparities.
Several studies further explore bias in retrieval-augmented systems, specifically within healthcare contexts. Levra et al.21 emphasized the risks of demographic biases that can be inadvertently amplified through retrieval processes in medical QA systems. Similarly, Pfohl et al. 23 introduced EquityMedQA, an open-ended dataset explicitly designed to test demographic biases in medical QA models, demonstrating disparities arising due to sensitive attributes such as race and gender. Recent methods have proposed sophisticated bias mitigation strategies to address these fairness issues. Chain-of-Thought (COT) prompting 24, 31 encourages explicit reasoning steps, potentially reducing reliance on demographic stereotypes. Counterfactual filtering 32 alters sensitive attributes to check model consistency and minimize discriminatory outcomes. Additionally, adversarial filtering and Majority Vote aggregation 33 have effectively reduced biases and promoted more equitable model outcomes by aggregating diverse demographic perspectives. Despite these advancements, challenges persist. Retrieval methods still sometimes select outdated or irrelevant information, as evidenced by error cases involving intersectional demographics (e.g., Asian non-binary individuals), identifying a gap in robustness and generalization capabilities. Ongoing work thus emphasizes the importance of developing trustworthy RAG systems, with comprehensive frameworks proposed to address reliability, privacy, and fairness comprehensively 22.
@@ Methods
We select four representative general-purpose large language models (LLMs) for integration into our Retrieval-Augmented Generation (RAG) framework, including DeepSeek-R1-8B , DeepSeek-R1-70B34, Meta-Llama-3-8B35, and PMC-LLaMA-13B 36. These models were chosen based on their proven capabilities in general reasoning, instruction-following tasks, and suitability for adaptation into domain-specific contexts, such as medical question-answering.
We apply four strategies to mitigate demographic bias in Retrieval-Augmented Generation (RAG). First, a Plain (baseline) condition involves no explicit intervention. Second, Chain-of-Thought (COT) Filtering 24, 31 guides models to produce step-by-step reasoning that isolates medical facts from demographic descriptions, aiming to lessen undesired influences of attributes like race or gender. Third, Counterfactual Filtering 32 prompts the model with methodically altered demographic labels and checks for inconsistencies or discriminatory behaviors in the resulting answers. Lastly, Adversarial Prompt Filtering reformulates queries to minimize reliance on socially sensitive markers, thus preventing the model from overfitting to or amplifying potential biases. We compare each filtering method’s output regarding accuracy, retrieval patterns, and fairness measures during inference.
@@
To further combat biases that may persist for a single demographic instance, we incorporate a majority voting approach 33 across multiple variants of the same question. Concretely, for each question in MedMCQA, MedQA, MMLU, and EquityMedQA, we generate distinct demographic variants by substituting race (Caucasian, African American, Asian, Hispanic) and gender (male, female, non-binary) into the query stem. For example, an original question “Which medication is recommended for a patient with chest pain?” can yield up to 12 variants if we combine four race attributes and three gender attributes. In practice, if both race and gender are not always simultaneously varied, we produce 4 to 6 variants (e.g., only race or only gender) depending on the experiment.
@@
(2) ROUGE-L. For open-ended QA such as EquityMedQA 23, which lacks fixed answer options, we measure correctness by comparing the generated text to a reference solution using ROUGE-L. This metric quantifies the overlap of the longest common subsequence (LCS) between the model’s output and the reference. A higher ROUGE-L indicates greater alignment with the reference’s content, helping detect factual completeness in a free-text generation.
@@
We adopt four corpora in our retrieval pipeline, each chunked into short snippets: (1) PubMed 37 for biomedical abstracts,(2) Medical Textbooks 26 for domain-specific knowledge, and (3) Wikipedia 38 for more general context. In addition, we combine these sources into a larger MedCorp if cross-domain retrieval is desired. Each snippet is indexed and retrieved via different retriever types, including a lexical approach (BM25) 39 and semantic encoders (Contriever) 40, 41. By default, we retrieve $k=15$ snippets per query; if multiple retrievers are used, we employ Reciprocal Rank Fusion (RRF) 42 to merge results.